OpenStudy (turingtest):

Quicky: Write\[\sin x-\cos x\]in the form\[A\sin(x+c)\]I totally forgot how to do this and can't find a reference.

OpenStudy (turingtest):

@satellite73 I know you can do this, please lend a hand.

OpenStudy (anonymous):

yes we can do it

OpenStudy (phi):

use sin(a+b)= sin(a)cos(b) + sin(b)cos(a)

OpenStudy (turingtest):

I just forgot how I used to know once...

OpenStudy (phi):

or in your case use sin(a-b)= sin(a)cos(b) - sin(b)cos(a)

OpenStudy (anonymous):

\[a\cos(\theta)+b\sin(\theta)=\sqrt{a^2+b^2}\sin(x+\tan^{-1}\frac{a}{b})\] if i am not mistaken

OpenStudy (anonymous):

we can derive the identity by using the addition angle formula

OpenStudy (turingtest):

so C is 3pi/4 ?

OpenStudy (anonymous):

in this case you will have a = -1, b = 1, \[\sqrt{2}\sin(x+\tan^{-1}(-1))=\sqrt{2}\sin(x-\frac{\pi}{4})\]

OpenStudy (turingtest):

...or - pi/4 okay thank you!

OpenStudy (anonymous):

yw now i will see if i can remember how to derive the identity, it is not hard

sam (.sam.):

to derive not hard at all

OpenStudy (turingtest):

I'd like to see it, but I may not stick around for the show getting ready for class

OpenStudy (turingtest):

what's up with the bike theme?

OpenStudy (turingtest):

satellite wrote\[a\cos(\theta)+b\sin(\theta)=\sqrt{a^2+b^2}\sin(x+\tan^{-1}\frac{a}{b})\]I know it doesn't matter in this case, but shouldn't it be\[a\sin(\theta)+b\cos(\theta)=\sqrt{a^2+b^2}\sin(x+\tan^{-1}\frac{a}{b})\]? (sin and cos switched, so the coefficient of sin is in the numerator when you take arctan)

OpenStudy (anonymous):

i am fairly sure i am right, that you should have \[\tan^{-1}(\frac{a}{b})\] where a is the coefficient of the cosine term it would be clearer if we derived the fromula

OpenStudy (turingtest):

okay, then I guess I will make the derivation a priority

OpenStudy (anonymous):

give me a second and i will start

OpenStudy (turingtest):

no rush, I may not be able to stick around but I'll read it when I return if nothing else

OpenStudy (anonymous):

ok i just need to refresh my memory.

OpenStudy (anonymous):

ok it is not hard

OpenStudy (anonymous):

put \[a\sin(x)+b\cos(x)=\frac{ra}{r}\sin(x)+\frac{br}{r}\cos(x)=r\left(\frac{a}{r}\sin(x)+\frac{b}{r}\cos(x)\right)\] where \[r=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

then note that we can find \[\alpha\] with \[\sin(\alpha)=\frac{b}{\sqrt{a^2+b^2}}, \cos(\alpha)=\frac{a}{\sqrt{a^2+b^2}}\] and now you will see why it really is the "a" on top and the "b" on the bottom

OpenStudy (phi):

Are you going to finish the derivation, or leave it to the reader?

OpenStudy (turingtest):

gotchya... just to be clear, you meant that is why in this case the "b" is on top (you put b as the coef of cos this time)

OpenStudy (anonymous):

\[a\sin(x)+b\cos(x)=r\left(\frac{a}{r}\sin(x)+\frac{b}{r}\cos(x)\right)\] \[=r\left(\cos(\alpha)\sin(x)+\sin(\alpha)\cos(x)\right)\] \[=r\sin(x+\alpha)\]

OpenStudy (turingtest):

so\[\alpha=\tan^{-1}\frac ba\]

OpenStudy (anonymous):

ok so lets see what we have, because i could have been wrong if it is \[a\sin(x)+b\cos(x)\] then you want \[\alpha\] with \[\cos(\alpha)=\frac{a}{r}\] \[\sin(\alpha)=\frac{b}{r}\] or \[\tan(\alpha)=\frac{b}{a}\] so \[\alpha =\tan^{-1}(\frac{b}{a})\] whew!

OpenStudy (turingtest):

got it, thanks sat! engraving this into my brain right now....